OpenStudy (anonymous):

I need help finishing this problem. I have it started but am stuck. Write a slope-intercept equation for a line w/ given characteristics: Passes through (4,0) and (-9,-3)

OpenStudy (anonymous):

Since we are given that the line passes through (4,0) and (-9,-3), the slope m=(-3-0)/(-9-4)=3/13. This ratio must hold for all two arbitrary points on this line. So let (x,y) be on the line. Then m=3/13=(y-0)/(x+3). From this we obtain y=(3/13)x-(12/13). You can check this by choosing (4,0) and (-9,-3) for (x,y).